# Statistical Taylor Expansion: A New and Path-Independent Method for Uncertainty Analysis 111

### v19

@@ 5 Polynomial
Because the maximal expansion term using Formula (5.1) is $\tilde{x}^{2N}$, $N$ in Formula (5.1) can reach only half of the maximal expansion order of Formula (2.6), for example, $224=448/2$ when the input uncertainty is assumed to be Gaussian, and $\hat{\kappa}=5$.
@@ 6 Matrix Calculations
$=\sum_{m=1\dots n}\sum_{<i_{1}\dots i_{m}>_{n}}\sum_{[j_{1}\dots j_{m}]_{n}}\mathbf{M}_{<i_{1}\dots i_{m}>_{n},[j_{1}\dots j_{m}]_{n}}\prod_{i=1\dots m}^{i\in\{i_{1}\dots i_{m}\}}\tilde{x}_{i,p_{i}};$ (6.2)
@@
$\delta^{2}|\mathbf{M}|$ $=\sum_{m=1}^{n}\sum_{<i_{1}\dots i_{m}>_{n}}\sum_{[j_{1}\dots j_{m}]_{n}}|\mathbf{M}_{<i_{1}\dots i_{m}>_{n},[j_{1}\dots j_{m}]_{n}}|^{2}\prod_{k=1\dots n}^{i_{k}\in\{i_{1}\dots i_{m}\}}(\delta x_{i_{k},j_{k}})^{2};$ (6.4)
Let $j_{1},j_{2}\dots j_{m}]_{n}$ denote a permutation of length $m$ from the vector $(1,2\dots n)$, and let $j_{1},j_{2}\dots j_{m}>_{n}$ for the ordered permutation of length $m$ [20]. Let $\mathbf{M}$ be a square matrix of size $n$ with element $x_{i,j}\pm\delta x_{i,j},i,j=1,2,\dots n$ at row index $i$ and column index $j$. Fourmulas (6.1) and (6.2) present the Taylor expansion of the determinant for matrix $\mathbf{M}$ when the uncertainties of matrix elements are all independent of each other, whereas Formulas (6.3) and (6.4) gives the statistical Taylor expansion of the determinant $|\mathbf{M}|$. In Formula (6.4), $\mathbf{M}_{<i_{1}\dots i_{m}>_{n},[j_{1}\dots j_{m}]_{n}}$ is a sub-matrix for $\mathbf{M}$, in which $i_{1}\dots i_{m}>_{n}$ contains the row indices, and $j_{1}\dots j_{m}]_{n}$ contains the column indices [20].
@@ 7 Mathematical Library Functions
Range $x\in[-100,+100]$ $x\in[1/32,32]$ $c\in[-3,+3]$ $x\in[-\pi,+\pi]$
Uncertainty $\delta x\in[10^{-15},1]$ $\delta x\in[10^{-15},0.2x]$ $\delta x\in[10^{-15},0.2]$ $\delta x\in[10^{-15},1]$
Error Deviation $1.000\pm 0.010$ $0.999\pm 0.011$ $0.989\pm 0.104$ $0.978\pm 0.150$
Table 1: The result error deviations of selected basic functions with Gaussian input noise $x\pm\delta x$ when $\delta x>10^{-15}$, obtained by variance arithmetic. The error deviation for $(1\pm\delta x)^{c}$ can be improved to $0.998\pm 0.034$ if diverging regions are excluded. The error deviation for $\sin(x\pm\delta x)$ can be improved to $1.000\pm 0.010$ if pole regions are excluded.
@@
Figure 17 shows that the error deviation for $\sin(x+\delta x)$ is $1.000\pm 0.010$, except approaching 0 when $x=\pm\pi/2$ and $\delta x<10^{-8}$. Near a distributional pole, the input uncertainty is suppressed, resulting in zero error deviation. As expected, $\delta^{2}\sin(x)$ exhibits the same periodicity as $\sin(x)$. The numerical errors of the library functions $\sin(x)$ and $\cos(x)$ over a larger range of $x$ are examined in greater detail in Section 9.
@@ 8.1 Moving-Window Linear Regression Algorithm
Formulas (8.1) and (8.2) provide the least-square line-fit of $Y=\alpha+\beta X$ between two set of data $Y_{j}$} and $X_{j}$}, where $j$ is an integer index identifying $(X,Y)$ pairs in the sets [5].
$\alpha$ $=\frac{\sum_{j}Y_{j}}{\sum_{j}1};$ (8.1)
$\beta$ $=\frac{\sum_{j}X_{j}Y_{j}\;\sum_{j}1-\sum_{j}X_{j}\;\sum_{j}Y_{j}}{\sum_{j}X_{j}X_{j}\;\sum_{j}1-\sum_{j}X_{j}\;\sum_{j}X_{j}};$ (8.2)
In many applications, data set $Y_{j}$} denotes an input data stream where $j$ represents the time index or sequence index. $Y_{j}$} is thus referred to as a time-series input, with $j$ corresponding to $X_{j}$. A moving window algorithm [5] is applied within a small window centered on each $j$. For each calculation window, $X_{j}=-H,-H+1\dots H-1,H$} where $H$ is an integer constant specifying the half width of the window. This choice ensures $\sum_{j}X_{j}=0$, which simplifies Formulas (8.1) and (8.2) into Formulas (8.3) and (8.4), respectively [17].
@@ 9.1 DFT (Discrete Fourier Transformation)
For example, the FT spectrum of a sine function is a Delta function at the signal frequency $f$ with a phase $\pi/2$ [5]. Figure 22 shows the DFT spectra of the sine function $h[k]=\sin(f\frac{2\pi}{128}k),k\in[0,127]$, where $f$ is its signal frequency. If DFT is regarded as the digital implementation of FT, the spectra exhibit no modeling error only when the input signal frequency $f$ is an integer, and display varying degrees of modeling errors otherwise. Because of these modeling errors, the use of DFT as the digital implementation of FT is questionable, even though such usage is ubiquitous, and fundamental to many areas of applied mathematics
@@ 9.2 FFT (Fast Fourier Transformation)
  • The forward and reverse transformations are identical except for a sign difference, implying that they are essentially the same algorithm, and any observed difference arises solely from the input data.
@@ 9.3 Testing Signals
For Sin or Cos signals, forward and reverse DFT transforms differ in their data prospective of time domain versus frequency domain:
  • The forward transformation converts a time-domain sine or cosine signal into a frequency-domain spectrum in which most values are zero, causing its uncertainties to grow more rapidly during cancellation.
  • In contrast, the reverse transformation spreads the precise frequency-domain spectrum (where most values are zero) back into a time-domain sine or cosine signal, causing its uncertainties to grow more slowly.
@@ 9.4 Trigonometric Library Errors
    The values of $\sin(\pi j/2^{L}),j\in[0,2^{L-2}]$ are the library sine directly, whereas the values of $\sin(\pi j/2^{L}),j\in[2^{L-2},2^{L-1}]$ are computed from the library $\cos(\pi(2^{L-1}-j)/2^{L})$.
@@ 9.5 Using Quart Sine for Sin/Cos Signals
Figure 25: Error distributions of normalized errors of Sin/Cos signals for forward, reverse and round-trip transformations (legend) using the Quart sine function. The FFT order is $18$.
@@ 9.6 Using Library Sine for Sin/Cos Signals
Figure 26: Error distributions of normalized errors of Sin/Cos signal for forward, reverse and round-trip transformations (legend) computed using the Library sine function. The FFT order is $18$.
@@
Figure 29: FFT value error waveform of $\sin(3\frac{2\pi}{2^{6}}j)$ computed using either the Library sine function or SciPy after the round-trip transformation. The legend distinguishes between the uncertainty and the value error. The x-axis represents the frequency index, and the y-axis represents both uncertainty and value error.
@@ 9.7 Using Quart Sine for Linear Signals
Figures 31 and 32 show the error deviations for the forward and the reverse transformations, respectively. The forward transformation exhibits a larger ideal coverage area than the reverse transformations: $\delta x>10^{-12}$ for the forward transformation, and $\delta x>10^{-8}$ for the reverse transformation. In other areas, both transformations achieve proper coverage with error values around $1$.
@@ 9.8 Using Library Sine for Linear Signals
Figures 35 and 36 show much smaller ideal coverage areas than those shown in Figures 31 and 32. Because uncertainty deviations grow more slowly in the reverse transformation than in the forward transformation, the former exhibits a smaller ideal coverage region. Outside of the ideal coverage region, proper coverage cannot be achieved for the reverse transformation. Furthermore, the range of input noise that produces ideal coverage decreases with increasing FFT order. At sufficiently high FFT orders (visually beyond FFT order $25$ for the reverse transformation), ideal coverage may no longer be achievable. Although FFT is widely regarded as one of the most robust numerical algorithms [5][10], and is generally insensitive to input errors, it can still fail because heavy calculation can amplify numerical errors in the Library sine function to completely obscure true signals. Such deterioration in calculation is not easily detectable when using conventional floating-point arithmetic.
Figure 37 shows that, even when variance arithmetic can no longer effectively track the value errors for either the forward or the reverse transformations, it can still effectively track the value errors for the round-trip transformation, as shown by the plateau region at high $L$ and low $\delta x$. Such error cancellation arises from dependence tracing in statistical Taylor expansion.
@@ 10 Regressive Generation of Sine
Formulas (10.2) and (10.3) calculate $\sin(\pi j/2^{L}),\cos(\pi j/2^{L}),j=0\dots 2^{L-2}$ repressively for regression order $L=0\dots 17$ starting from Formula (10.1). Formula (10.4) shows that such regression guarantees both $\sin(x)^{2}+\cos(x)^{2}=1$ and $\sin(2x)=2\sin(x)\cos(x)$, such that value errors will not accumulate when the regression order $L$ increases.
@@ 11.2 Improvements Needed
Bound momentum $\zeta(n,\kappa)$ should be extended to all probability distributions. The choice of ideal bounding range $\hat{\kappa}$ should be extended to other distributions, perhaps by using the same value of linear leakage $\varepsilon\equiv 1-\zeta(2,\kappa)$. The procedure to find the bound moments $\kappa$ for each sample count $N$ should be developed for discrete distributions. The statistical meaning for ideal leakage $\varepsilon\in[0,1]$ in the context of its choice of ideal bounding range $\hat{\kappa}$ needs further investigation: Is $\varepsilon$ the possibility for the result to be not bound by $\hat{\kappa}$?
@@
As an enhancement to dependency tracing, source tracing identifies each input’s contribution to the overall result uncertainty. This capability enables engineers to pinpoint the primary sources of measurement inaccuracy and guides them to targeted improvements in data acquisition and processing strategies. For example, Formula (2.33) can provide guidance on improving the ideal leakage of $x\pm y$.
@@ 11.3 Possible Connections to Quantum Physics
For input distribution without bounding, Formula (2.6) may not converge at all, for example, when the input uncertainty is Laplace [4], $\zeta(2n,\infty)=(2n)!$. Formula (2.6) converges in all cases only when the input uncertainty is bounded because of sampling, which resembles the need for re-normalization in quantum field theory [21]. Furthermore, the quantitative dependence of convergence on the choice of ideal bounding range $\hat{\kappa}$ is similar to the principle of quantum physics: What it is depends on how it is measured. For example, it is possible that an experimental measurement reduces $\hat{\kappa}$ for a wave function, to cause the wave function to collapse. Both resemblances are worthy of further investigation.

### v20

@@ 5 Polynomial
Because the maximal expansion term using Formula (5.1) is $\tilde{x}^{2N}$, $N$ in Formula (5.1) can reach only half of the maximal expansion order of Formula (2.6), for example, $2j=224=448/2$ for $\zeta(2n,5)$ of Gaussian input uncertainty.
@@ 6 Matrix Calculations
$\delta^{2}|\mathbf{M}|$ $=\sum_{m=1}^{n}\sum_{<i_{1}\dots i_{m}>_{n}}\sum_{[j_{1}\dots j_{m}]_{n}}|\mathbf{M}_{<i_{1}\dots i_{m}>_{n},<[j_{1}\dots j_{m}]>_{n}}|^{2}\prod_{k=1\dots m}\zeta_{i,j}(2,\kappa_{i,j})(\delta x_{i_{k},j_{k}})^{2};$ (6.3)
Let $j_{1},j_{2}\dots j_{m}>_{n}$ denote a selection of $m$ numbers from the sequence of natural numbers $1,2\dots n$ in sequence, which is named as the ordered permutation. Let $j_{1},j_{2}\dots j_{m}]_{n}$ and $(-1)^{[j_{1},j_{2}\dots j_{m}]_{n}}$ denote a permutation of $j_{1},j_{2}\dots j_{m}]>_{n}$ and the corresponding sign [21], in which $j_{1},j_{2}\dots j_{m}]>_{n}$ denote the corresponding ordered permutation. Let $\mathbf{M}$ be a square matrix of size $n$ with element $x_{i,j}\pm\delta x_{i,j},i,j=1,2\dots n$ at row index $i$ and column index $j$. Fourmula (6.1) presents the Taylor expansion of the determinant for matrix $\mathbf{M}$ when the uncertainties of matrix elements are all independent of each other, whereas Formulas (6.2) and (6.3) gives the statistical Taylor expansion of the determinant $|\mathbf{M}|$. In Formula (6.3), $\mathbf{M}_{<i_{1}\dots i_{m}>_{n},[j_{1}\dots j_{m}]_{n}}$ is a sub-matrix for $\mathbf{M}$, in which $i_{1}\dots i_{m}>_{n}$ contains the row indices, and $j_{1}\dots j_{m}]_{n}$ contains the column indices [21].
@@ 7 Mathematical Library Functions
Range $x\in[-100,+100]$ $x\in[1/32,32]$ $c\in[-3,+3]$ $x\in[-\pi,+\pi]$
Uncertainty $\delta x\in[10^{-15},1]$ $P(x)\in[10^{-15}x,0.2x]$ $P(x)\in[10^{-15},0.2]$ $\delta x\in[10^{-15},1]$
Error Deviation $1.000\pm 0.010$ $0.999\pm 0.011$ $0.989\pm 0.104$ $0.997\pm 0.027$
Table 1: The result error deviations of selected basic functions with Gaussian input noise $x\pm\delta x$ when $\delta x>10^{-15}$, obtained by variance arithmetic. The error deviation for $(1\pm\delta x)^{c}$ can be improved to $1.000\pm 0.007$ if diverging regions are excluded. The error deviation for $\sin(x\pm\delta x)$ can be improved to $1.000\pm 0.010$ if pole regions are excluded.
@@
Figure 17 shows that the error deviation for $\sin(x+\delta x)$ is $1.000\pm 0.010$, except approaching 0 when $x=\pm\pi/2$ and $\delta x<10^{-8}$. Near a distributional pole, the input uncertainty is suppressed, resulting in zero error deviation. The numerical errors of the library functions $\sin(x)$ and $\cos(x)$ over a larger range of $x$ are examined in greater detail in Section 9.
@@ 8.1 Moving-Window Linear Regression Algorithm
In many applications, data set $Y_{j}$} denotes an input data stream where $j$ represents the time index or sequence index. $Y_{j}$} is thus referred to as a time-series input. A moving window algorithm [5] is applied within a small window centered on each $j$. For each calculation window, Formulas (8.1) and (8.2) present the least-square line-fit of $Y=\alpha+\beta j$ [17].
@@ 9.1 DFT (Discrete Fourier Transformation)
However, DFT if not a faithful digital implementation of continuous Fourier transform (FT) [5]. For example, the FT spectrum of a sine function is a Delta function at the signal frequency $f$ with a phase $\pi/2$ [5]. Figure 22 shows the DFT spectra of the sine function $h[k]=\sin(f\frac{2\pi}{128}k),k\in[0,127]$, where $f$ is its signal frequency. The DFT spectra exhibit no modeling error only when the input signal frequency $f$ is an integer, and display varying degrees of modeling errors otherwise. Because of these modeling errors, the use of DFT as the digital implementation of FT is questionable, even though such usage is ubiquitous, and fundamental to many areas of applied mathematics [5].
@@ 9.3 Testing Signals
The forward and reverse transformations are identical except for a sign difference, implying that they are essentially the same algorithm, and any observed difference arises mainly from the input data. For Sin or Cos signals:
  • The forward transformation converts a time-domain sine or cosine signal into a frequency-domain spectrum in which most values are zero, causing its uncertainties to grow more rapidly during mutual cancellation of input data.
  • The reverse transformation spreads the precise frequency-domain spectrum (where most values are zero) back into a time-domain sine or cosine signal, causing its uncertainties to grow more slowly.
@@ 9.4 Trigonometric Library Errors
    The values of $\sin(\pi j/2^{L}),j\in[0,2^{L-2}]$ are the library sine directly, whereas the values of $\sin(\pi j/2^{L}),j\in[2^{L-2},2^{L-1}]$ are computed from the library $\cos(\pi(2^{L-1}-j)/2^{L}),j\in[0,2^{L-2}]$.
@@ 9.5 Using Quart Sine for Sin/Cos Signals
Figure 25: Error distributions of normalized errors of Sin/Cos signals without added input noise for forward, reverse and round-trip transformations (legend) using the Quart sine function. The FFT order is $18$.
@@ 9.6 Using Library Sine for Sin/Cos Signals
Figure 26: Error distributions of normalized errors of Sin/Cos signal without added input noise for forward, reverse and round-trip transformations (legend) computed using the Library sine function. The FFT order is $18$.
@@
Figure 29: FFT value error waveform of $\sin(3\frac{2\pi}{2^{6}}j)$ computed using either the Library sine function or SciPy after the round-trip transformation. The legend distinguishes between the uncertainty and the value error. The x-axis represents the time index, and the y-axis represents both uncertainty and value error.
@@ 9.7 Using Quart Sine for Linear Signals
Figures 31 and 32 show the error deviations for the forward and the reverse transformations, respectively, both of which resemble Figure 14. The forward transformation exhibits a larger ideal coverage area than the reverse transformations: $\delta x>10^{-12}$ for the forward transformation, and $\delta x>10^{-8}$ for the reverse transformation. In other areas, both transformations achieve proper coverage with error values around $1$.
@@ 9.8 Using Library Sine for Linear Signals
Figures 35 and 36 show much smaller ideal coverage areas than those shown in Figures 31 and 32, respectively. Because uncertainty deviations grow more slowly in the reverse transformation than in the forward transformation, the former exhibits a smaller ideal coverage region. Outside of the ideal coverage region, proper coverage cannot be achieved for the reverse transformation. Furthermore, the range of input noise that produces ideal coverage decreases with increasing FFT order. At sufficiently high FFT orders (visually beyond FFT order $25$ for the reverse transformation), ideal coverage may no longer be achievable. Although FFT is widely regarded as one of the most robust numerical algorithms [5][10], and is generally insensitive to input errors, it can still fail because heavy calculation can amplify numerical errors in the Library sine function to completely obscure true signals. Such deterioration in calculation is not easily detectable when using conventional floating-point arithmetic.
Figure 37 shows that, even when variance arithmetic can no longer effectively track the value errors for either the forward or the reverse transformations, it can still effectively track the value errors for the round-trip transformation, as shown by the plateau region of error deviations at high $L$ and low $\delta x$. Such error cancellation arises from dependence tracing in statistical Taylor expansion.
@@ 10 Regressive Generation of Sine
Formulas (10.2) and (10.3) calculate $\sin(j\pi/2^{L}),\cos(j\pi/2^{L}),j=0\dots 2^{L-2}$ repressively for regression order $L=0\dots 17$ starting from Formula (10.1). Formula (10.4) shows that such regression guarantees both $\sin(x)^{2}+\cos(x)^{2}=1$ and $\sin(2x)=2\sin(x)\cos(x)$, such that value errors will not accumulate when the regression order $L$ increases.
@@ 11.2 Improvements Needed
Bound momentum $\zeta(n,\kappa)$ should be extended to all probability distributions. The choice of ideal bounding range $\hat{\kappa}$ should be extended to other distributions, perhaps by using the same value of linear leakage $\varepsilon\equiv 1-\zeta(2,\kappa)$. The procedure to find the bound moments $\kappa$ for each sample count $N$ should be developed for discrete distributions.
@@
As an enhancement to dependency tracing, source tracing identifies each input’s contribution to the overall result uncertainty. This capability enables engineers to pinpoint the primary sources of measurement inaccuracy and guides them to targeted improvements in data acquisition and processing strategies. For example, Formulas (2.30) and (2.31) can provide guidance on improving the ideal leakage of $x\pm y$ and $xy$ respectively.
@@ 11.3 Possible Connections to Quantum Physics
For input distribution without bounding, Formula (2.6) may not converge at all, for example, when the input uncertainty is Laplace [4], $\zeta(2n,\infty)=(2n)!$. Formula (2.6) converges in all cases only when the input uncertainty is bounded because of sampling, which resembles the need for re-normalization in quantum field theory [22]. Furthermore, the quantitative dependence of convergence on the choice of ideal bounding range $\hat{\kappa}$ is similar to the principle of quantum physics: What it is depends on how it is measured. For example, it is possible that an experimental measurement reduces $\hat{\kappa}$ for a wave function, to cause the wave function to collapse when the uncertainty deviation reduces from infinite to finite. Both resemblances are worthy of further investigation.